OpenStudy (anonymous):

1. A local movie theater charges $12 for an adult ticket and $10 for a child’s ticket. A group of eight people spent a total of $86 on tickets to a movie. How many adults and how many children were in the group? a. Write a system of linear equations based on the description. Use x to represent the number of adults and y to represent the number of children.

OpenStudy (johnweldon1993):

Well we know that they spent a total of 86 on tickets. 12 each for adults 10 each for children's so we have 12a + 10c = 86 *sorry letting x = adults and y = childrens 12x + 10y = 86

OpenStudy (anonymous):

ok thank you i just wanted to make sure my equation was right

OpenStudy (johnweldon1993):

Also have the other equation to do ... there were 8 total people so x + y = 8

OpenStudy (anonymous):

i have another question to if you would like to help?

OpenStudy (anonymous):

its the same problem to

OpenStudy (johnweldon1993):

yeah no problem

OpenStudy (anonymous):

Solve the system using the elimination method. Show all the steps leading to your answer

OpenStudy (johnweldon1993):

Oh, so now we solve the system using elimination...okay...well we had our 2 equations x + y = 8 12x + 10y = 86 We want to eliminate a variable...lets eliminate 'y' multiply that top equation by -10 -10x - 10y = -80 12x + 10y = 86 Now when we ADD the equations together...the -10y and +10y cancel out..so we have -10x + 12x = ...? -80 + 86 = ...?

OpenStudy (anonymous):

We add the top and bottom equations together?

OpenStudy (johnweldon1993):

Yes that is how the elimination method works...you make it so the 2 equations "when added together" get rid of 1 variable..making it possible to solve for the other variable

OpenStudy (anonymous):

-90 +98 =8 ? I dont know I'm not that smart when it comes to this.

OpenStudy (johnweldon1993):

Don't worry about it so when we have the 2 equations x + y = 8 12x + 10y = 86 We want to multiply the top equation...by some number to eliminate one of the variables...I chose to eliminate 'y' to get rid of 10y...we need to have a -10y..because -10y + 10y = 0 (canceled) so we multiply everything in that top equation by -10 -10x - 10y = -80 now the second equation stays the same 12x + 10y = 86 Now add them -10x + 12x = ...? -10y + 10y = ...? -80 + 86 = ...?

OpenStudy (anonymous):

Ok hold on this might take a moment

OpenStudy (johnweldon1993):

No problem!

OpenStudy (anonymous):

-10x + 12x = 2 -10y + 10y = 0 -80 + 86 = 6

OpenStudy (johnweldon1993):

Numbers = correct...but remember the variables stay when you add -10x + 12x = 2x (keep the variable) great...so now we have 2x = 6 what does 'x' equal ...?

OpenStudy (anonymous):

3?

OpenStudy (johnweldon1993):

Right...x = 3 This says that you had 3 adults that went Now we can go back to our original equations x + y = 8 we now know 'x'... = 3....so lets find 'y' 3 + y = 8 y = ...?

OpenStudy (anonymous):

5!

OpenStudy (johnweldon1993):

Right! So we have 3 adults...and 5 children Lets check it We had another equation right? 12x + 10y = 86 plug in x = 3 and y = 5 and see if it works...

OpenStudy (anonymous):

You want me to use it on that equation?.. how would i do that? 12 x 3 + 10 x 5 = 86?

OpenStudy (johnweldon1993):

Exactly...We already have the answer...this is just how you check it... we know 'x' = 3 and 'y' = 5 So we replace 'x' and 'y' with those values 12x + 10y = 86 12(3) + 10(5) = 86

OpenStudy (anonymous):

Ohhh i see so thats the answer?

OpenStudy (johnweldon1993):

12(3) + 10(5) = 86 36 + 50 = 86 86 = 86 So now we know for SURE that there were 3 adults and there were 5 children

OpenStudy (anonymous):

Ok thank you! i just need one more question to finish this problem. Write 1-2 sentences to explain what the solution means for this situation

OpenStudy (anonymous):

So i just write down the steps right?

OpenStudy (johnweldon1993):

Well yeah I suppose...you could say....the solution means that there were 3 adults and 5 children that went to the movies Something to that extent...

OpenStudy (anonymous):

Ok thank you for the help. i appreciate it man. I really needed to get this assignment done and you made it really simple.

OpenStudy (johnweldon1993):

No problem!